# Generation of complete graph states in a spin-$1/2$ Heisenberg chain with a globally optimized magnetic field

### 07 Jan 2024 v2

@@ IV.2 Experimental feasibility analysis
This part focuses on the experimental feasibility by examining the experimental imperfections and providing a concise design for an experiment procedure using the Rydberg atom array. This includes the production of the initial state and the final decoupling process.
@@ IV.2.1 The influence of experimental errors
Here, we rigorously assess the scheme by taking into account experimental imperfections, such as spontaneous radiation of Rydberg levels, as well as the discrepancy between the dipole-dipole interaction $V_{\textrm{dip}}$ and the magnetic field $B(t)$. The associated numerical simulations are conducted employing Gaussian-type optimal pulses.
@@
where $\gamma_{\uparrow(\downarrow)}$ is the decay rate of the Rydberg state |()$\uparrow$($\downarrow$)$\rangle$| ↑ ( ↓ ) ⟩ equals to $1/\tau_{\uparrow(\downarrow)}$, and $s_{i}=|g\rangle_{i}\langle j|$ denotes the decay channel of the $i$th atom. Through numerical simulation, dissipation is observed to lead to a reduction in the population of the target state by values of $\{0.0006,0.0009,0.0013,0.0017\}$, corresponding to $N$ ranging from three to six, respectively.
@@
Figure 4: The influence of the mismatch of the dipole interaction $V_{\textrm{dip}}$ and the magnetic field $B(t)$ caused by atomic vibration. (a) Populations of the complete multiparticle graph states with range error $\delta r$ under the Hamiltonian $H_{\textrm{sys}}(\delta r)$. (b) and (c) The average evolution results for $N=4,6$ by considering a true fluctuation where the position distribution is Gaussian with variance $\{\sigma_{x},\sigma_{y},\sigma_{z}\}\simeq\{193.5,193.5,1242.9\}~{}{\rm nm}$. Note that the light red region shows the results of 50 stochastic simulations, and the dashed lines correspond to the average results.
@@
$H_{\delta V}(t)=\sum_{\langle i,j\rangle}V_{\textrm{dip}}(\delta r)(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j})+\sum_{i=1}^{N}B(t)S^{z}_{i}.$ (16)
@@
It is worth noting that as the vdW interaction is also contingent on the distance between atoms, and the Hamiltonian $H_{\textrm{err}}$ should likewise be formulated as
$H^{\textrm{err}}_{\delta V}$ $=$ $\sum_{\langle i,j\rangle}[U^{\uparrow}_{ij}(\delta r)|\uparrow_{i}\uparrow_{j}\rangle\langle\uparrow_{i}\uparrow_{j}|+U^{\downarrow}_{ij}(\delta r)|\downarrow_{i}\downarrow_{j}\rangle\langle\downarrow_{i}\downarrow_{j}|]$ (17)
$+\sum_{\langle\langle i,j\rangle\rangle}\frac{V_{\textrm{dip}}(\delta r)}{8}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}),$
@@
where $U^{\uparrow(\downarrow)}(\delta r)=-C^{\uparrow(\downarrow)}_{6}/(R+\delta r)^{6}$. In Fig. 4(a), we present the evolution results governed by $H_{\textrm{sys}}(\delta r)=H_{\delta V}(t)+H^{\textrm{err}}_{\delta V}$, with $N$ varying from three to six, respectively. The population of target states $P(N)_{O}$ consistently exceeds 0.9 for $\delta r\leq 300~{}{\rm nm}$. In experimental settings, the vibration of atoms near the ideal position, induced by nonzero temperature, follows the Maxwell-Boltzmann distribution and is associated with the atomic temperature as well as the parameters of the tweezer beam and trap. According to Ref. Fu et al. (2022), the atomic temperature can be cooled to $5.2~{}\mu{\rm K}$ in a $50~{}\mu{\rm K}$ trap. Setting the corresponding parameters of the laser beams as wavelength $\lambda_{f}=830~{}{\rm nm}$, typical beam power $P_{f}=174~{}\mu{\rm W}$, and waist ($1/e^{2}$ intensity radius) $\omega_{f}=1.2~{}\mu{\rm m}$, we estimate a position distribution in Gaussian with variance $\{\sigma_{x},\sigma_{y},\sigma_{z}\}\simeq\{193.5,193.5,1242.9\}~{}{\rm nm}$. Thus, in Fig. 4(b) and Fig. 4(c), we further assess the robustness of the protocol with $N=4$ and $N=6$ as realistically as possible, considering a three-dimensional random vibration under these estimated parameters. The distance between the atoms $R$ is redefined as $D=|\textbf{R}_{i}-\textbf{R}_{i+1}|$, where $\textbf{R}_{i}=(x_{i},y_{i},z_{i})$. Since the evolution time $T$ is less than $0.25~{}\mu{\rm s}$, we reasonably consider only one group of random fluctuations ($\sigma_{x,y,z}$) error throughout the evolution process here. Fig. 4(b) and 4(c) respectively illustrate the population of the target states $|K_{4}\rangle$ and $|K_{6}\rangle$, where the light red portions represent the results of 50 stochastic simulations, and the dashed lines in dark blue correspond to the average outcomes. The average populations of $|K_{4}\rangle$ and $|K_{6}\rangle$ in this test are $0.9847$ and $0.9310$, respectively. With advancements in experimental methods, the actual vibration of atoms in the trap is expected to be smaller than our estimated parameters Labuhn et al. (2016); Chew et al. (2022), which is more conducive to the realization of our scheme.
@@
For the latter reason, the intensity fluctuation of the magnetic field is introduced by a time-dependent fluctuation $\delta B$. It is assumed to follow the normal distribution functions with the standard deviations $\sigma(t)\simeq 2\pi\times 0.5~{}{\rm MHz}$. Thus, we have the system Hamiltonian
@@ IV.2.2 Mapping between Rydberg states and ground states
(i) The preparation of the initial state. In this protocol, the initial state is selected as
@@
Fig. 7 illustrates a complete evolution with $N=3$. To enhance observability, we depict the evolution of the target states for each process, where $|\psi_{1}\rangle=\bigotimes_{i=1}^{3}[{1}/{\sqrt{2}}(|\uparrow_{i}\rangle-i|\downarrow_{i}\rangle)]$, $|\psi_{2}\rangle={1}/{(2\sqrt{2})}\bigotimes_{i=1}^{3}[|\uparrow_{i}\rangle(-1)^{3-i}\prod_{j=i+1}^{3}\sigma_{z}^{j}-i|\downarrow_{i}\rangle]$, $|\psi_{3}\rangle={1}/{(2\sqrt{2})}\bigotimes_{i=1}^{3}[|\uparrow_{i}\rangle(-1)^{3-i}\prod_{j=i+1}\sigma_{z}^{j}-|r_{i}\rangle]$, and $|\psi_{4}\rangle={1}/{(2\sqrt{2})}\bigotimes_{i=1}^{3}[-|0_{i}\rangle(-1)^{3-i}\prod_{j=i+1}\sigma_{z}^{j}+|1_{i}\rangle]$. The dashed and dotted lines correspond to the initial state preparation and the decoupling process, respectively, while the solid line represents the core evolution with the optimized $B(t)$ in Gaussian type. Under the parameters $\Omega=2\pi\times 4~{}{\rm MHz}$, $\Omega_{\rm mw}^{a}=2\pi\times 70~{}{\rm MHz}$ and $\Omega_{\textrm{mw}}^{b}=2\pi\times 200~{}{\rm MHz}$, the population of the target state can reach about $0.9916$ with a total evolution time $t_{\textrm{tot}}=0.3971~{}\mu{\rm s}$. Given that the independent evolution process of the core operation can reach perfection with $N=3$, the error caused by the preparation of the initial state and the return driving is approximately $0.0072$.
@@
Initial preparing and decoupling 0.0072
Spontaneous radiation 0.0006 0.0009 0.0013 0.0017
Atomic vibration (±$\pm 100~{}{\rm nm}$) 0.0008 0.0022 0.0041 0.0051
$P(N)_{O}$ 0.9902 0.9808 0.9655 0.9303
As shown in Table. 3, we make an estimation on the population of complete states with $N=3\sim 6$ considering the experimental errors to make a comparison with the other protocol. In recent experiments, the fidelity of the two-qubit entanglement gate has been improved to $99.5\%$ Jandura and Pupillo (2022); Evered et al. (2023). On the basis of this, we can estimate the time and fidelity required to achieve the complete graph states using CZ gates to make a comparison. For $N=3$, we need at least three CZ gates, and the estimated fidelity of the complete graph state is about $0.985$, which is slightly lower than our protocol. For $N=6$, fifteen CZ gates are needed leading to the estimated fidelity $0.928$ which is on par with our scheme. However, such CZ gate operations cannot be performed simultaneously, but step by step. Although a single gate operation time may be less than $300~{}{\rm ns}$, the evolution time required to prepare the target state is difficult to estimate to be less than $0.25~{}\mu{\rm s}$. It increases rapidly with the increase of the number of particles $N$. In contrast, our solution greatly reduces the time required to prepare multiparticle complete graph states while maintaining high fidelity. This reduction is crucial from the perspective of decoherence.
@@
In summary, we have introduced an efficient single-step technique for generating multiparticle complete graph states. Our approach relies solely on a constant nearest-neighbor interaction strength and a time-varying magnetic field which is optimized using the GRAPE algorithm. This methodology is applicable to various physical systems capable of realizing a spin-$1/2$ Heisenberg $XX$ chain, such as superconducting systems, trapped ion systems, and neutral-atom systems. We have thoroughly investigated the protocol’s execution in the Rydberg atom array system, evaluating its resilience in the face of experimental errors. Our numerical findings underscore the robustness of our approach against challenges like atomic position inaccuracies and pulse oscillations. Our one-step solution not only significantly reduces the operating time, but also maintains a high estimated fidelity, presenting a notable departure from quantum gate-based methods traditionally employed for creating graph states. We anticipate that this discovery will quickly contribute to the practical implementation of quantum computing and quantum error correction in neutral-atom systems.
Figure 8: The evolution governed by the Hamiltonian (20) with $N=3$. (a) The intensity of the constant-amplitude pulse $B$ satisfies the relationship shown in Eq. (24). (b) The scanning results of the target state population with different parameters.
@@ acknowledgment
The authors thank Professor Re-Bing Wu of Tsinghua University for his guidance on the GRAPE optimal algorithm. The anonymous reviewers are also thanked for constructive comments that helped to improve the quality of this paper. This work is supported by the NFSC through Grant No. 12174048, the Fundamental Research Funds for the Central Universities (Grant No. 3122023QD26), and the Research Start-up Funds of the Shenyang Normal University (Grants No. BS202301, No. BS202303, No. BS202314).
@@ Appendix A A protocol for $N=3$ under constant amplitude magnetic field
$\psi_{1}(t)=\frac{e^{-\frac{3itB}{2}}}{2\sqrt{2}},$
$\psi_{2}(t)=\frac{e^{-i\frac{1}{2}t(4\sqrt{2}J+B)}[2+\sqrt{2}-(\sqrt{2}-2)e^{4i\sqrt{2}Jt}]}{8\sqrt{2}},$
$\psi_{3}(t)=\frac{e^{-i\frac{1}{2}t(4\sqrt{2}J+B)}[1+\sqrt{2}-(\sqrt{2}-1)e^{4i\sqrt{2}Jt}]}{4\sqrt{2}},$
$\psi_{4}(t)=\frac{1}{4}e^{\frac{itB}{2}}(\sqrt{2}\cos{\frac{4Jt}{\sqrt{2}}}-i\sin{\frac{4Jt}{\sqrt{2}}}),$
$\psi_{5}(t)=\frac{e^{-i\frac{1}{2}t(4\sqrt{2}J+B)}[2+\sqrt{2}-(\sqrt{2}-2)e^{4i\sqrt{2}Jt}]}{8\sqrt{2}},$
$\psi_{6}(t)=\frac{1}{4}e^{\frac{itB}{2}}(\sqrt{2}\cos{\frac{4Jt}{\sqrt{2}}}-2i\sin{\frac{4Jt}{\sqrt{2}}}),$
$\psi_{7}(t)=\frac{1}{4}e^{\frac{itB}{2}}(\sqrt{2}\cos{\frac{4Jt}{\sqrt{2}}}-i\sin{\frac{4Jt}{\sqrt{2}}}),$

### 11 Mar 2024 v3

@@ IV.2 Experimental feasibility analysis
This part focuses on experimental feasibility by examining the experimental imperfections and providing a concise design for an experiment procedure using the Rydberg atom array. This includes the production of the initial state and the final decoupling process.
@@ IV.2.1 The influence of experimental errors
Here, we rigorously assess the scheme by taking into account experimental imperfections, such as spontaneous radiation of Rydberg levels, as well as the discrepancy between the dipole-dipole interaction $V_{\textrm{dip}}$ and the magnetic field $B(t)$. The associated numerical simulations employ Gaussian-type optimal pulses.
@@
where $\gamma_{\uparrow(\downarrow)}$ is the decay rate of the Rydberg state |()$\uparrow$($\downarrow$)$\rangle$| ↑ ( ↓ ) ⟩ equals to $1/\tau_{\uparrow(\downarrow)}$, and $s_{i}=|g\rangle_{i}\langle j|$ denotes the decay channel of the $i$th atom. Through numerical simulation, dissipation is observed to lead to a reduction in the population of the target state by values of $\{0.0006,0.0009,0.0010,0.0017\}$, corresponding to $N$ ranging from three to six, respectively.
@@
Figure 4: The influence of the mismatch of the dipole interaction $V_{\textrm{dip}}$ and the magnetic field $B(t)$ caused by atomic vibration. (a) Populations of the complete multiparticle graph states with range error $\delta r$ under the Hamiltonian $H_{\textrm{sys}}(\delta r)$. (b) and (c) Average evolution results for $N=4,6$ considering a true fluctuation where the position distribution is Gaussian with variance $\{\sigma_{x},\sigma_{y},\sigma_{z}\}\simeq\{193.5,193.5,1242.9\}~{}{\rm nm}$. Note that the light-red region shows the results of 50 stochastic simulations, and the dashed lines correspond to the average results.
@@
$H_{\delta V}(t)=\sum_{\langle i,j\rangle}\frac{V_{\textrm{dip}}(\delta r)}{2}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j})+\sum_{i=1}^{N}B(t)S^{z}_{i}.$ (16)
@@
It should be noted that as the vdW interaction also depends on the distance between the atoms, the Hamiltonian $H_{\textrm{err}}$ should likewise be formulated as
$H^{\textrm{err}}_{\delta V}$ $=$ $\sum_{j>i}[U^{\uparrow}_{ij}(\delta r)|\uparrow_{i}\uparrow_{j}\rangle\langle\uparrow_{i}\uparrow_{j}|+U^{\downarrow}_{ij}(\delta r)|\downarrow_{i}\downarrow_{j}\rangle\langle\downarrow_{i}\downarrow_{j}|]$ (17)
$+\sum_{j>i+1}\frac{V_{\textrm{dip}}^{ij}(\delta r)}{2}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}),$
@@
where $U^{\uparrow(\downarrow)}(\delta r)=-C^{\uparrow(\downarrow)}_{6}/(R_{ij}+\delta r)^{6}$, and $V_{\rm dip}^{ij}=C_{3}(1-3\cos^{2}\theta)/(R_{ij}+\delta r)^{3}$. In Fig. 4(a), we present the evolution results governed by $H_{\textrm{sys}}(\delta r)=H_{\delta V}(t)+H^{\textrm{err}}_{\delta V}$, with $N$ varying from three to six, respectively. The population of target states $P(N)_{O}$ consistently exceeds 0.9 for $\delta r\leq 300~{}{\rm nm}$. In experimental settings, the vibration of atoms near the ideal position, induced by nonzero temperature, follows the Maxwell-Boltzmann distribution and is associated with the atomic temperature as well as the parameters of the tweezer beam and trap. According to Ref. [71], the atomic temperature can be cooled to $5.2~{}\mu{\rm K}$ in a $50~{}\mu{\rm K}$ trap. Setting the corresponding parameters of the laser beams as wavelength $\lambda_{f}=830~{}{\rm nm}$, typical beam power $P_{f}=174~{}\mu{\rm W}$, and waist ($1/e^{2}$ intensity radius) $\omega_{f}=1.2~{}\mu{\rm m}$, we estimate a position distribution in Gaussian with variance $\{\sigma_{x},\sigma_{y},\sigma_{z}\}\simeq\{193.5,193.5,1242.9\}~{}{\rm nm}$. Thus, in Fig. 4(b) and Fig. 4(c), we further assess the robustness of the protocol with $N=4$ and $N=6$ as realistically as possible, considering a three-dimensional random vibration under these estimated parameters. The distance between the atoms $R$ is redefined as $D=|\textbf{R}_{i}-\textbf{R}_{i+1}|$, where $\textbf{R}_{i}=(x_{i},y_{i},z_{i})$. Since the evolution time $T$ is less than $0.25~{}\mu{\rm s}$, we reasonably consider only one group of random fluctuations ($\sigma_{x,y,z}$) error throughout the evolution process here. Fig. 4(b) and 4(c) respectively illustrate the population of the target states $|K_{4}\rangle$ and $|K_{6}\rangle$, where the light red portions represent the results of 50 stochastic simulations, and the dashed lines in dark blue correspond to the average outcomes. The average populations of $|K_{4}\rangle$ and $|K_{6}\rangle$ in this test are $0.9728$ and $0.9187$, respectively. With advancements in experimental methods, the actual vibration of the atoms in the trap is expected to be smaller than our estimated parameters [72, 73], which is more conducive to the realization of our scheme.
@@
For the latter reason, the fluctuation of the magnetic field intensity is introduced by a time-dependent fluctuation $\delta B$. It is assumed to follow normal distribution functions with standard deviations $\sigma(t)\simeq 2\pi\times 0.5~{}{\rm MHz}$. Thus, we have the system Hamiltonian
@@ IV.2.2 Mapping between Rydberg states and ground states
(i) Preparation of the initial state. In this protocol, the initial state is selected as
@@
Fig. 7 illustrates a complete evolution with $N=3$. To enhance observability, we depict the evolution of the target states for each process, where $|\psi_{1}\rangle=\bigotimes_{i=1}^{3}[{1}/{\sqrt{2}}(|\uparrow_{i}\rangle-i|\downarrow_{i}\rangle)]$, $|\psi_{2}\rangle={1}/{(2\sqrt{2})}\bigotimes_{i=1}^{3}[|\uparrow_{i}\rangle(-1)^{3-i}\prod_{j=i+1}^{3}\sigma_{z}^{j}-i|\downarrow_{i}\rangle]$, $|\psi_{3}\rangle={1}/{(2\sqrt{2})}\bigotimes_{i=1}^{3}[|\uparrow_{i}\rangle(-1)^{3-i}\prod_{j=i+1}\sigma_{z}^{j}-|r_{i}\rangle]$, and $|\psi_{4}\rangle={1}/{(2\sqrt{2})}\bigotimes_{i=1}^{3}[-|0_{i}\rangle(-1)^{3-i}\prod_{j=i+1}\sigma_{z}^{j}+|1_{i}\rangle]$. The dashed and dotted lines correspond to the initial state preparation and the decoupling process, respectively, while the solid line represents the core evolution with the optimized $B(t)$ in Gaussian type. Under the parameters $\Omega=2\pi\times 4~{}{\rm MHz}$, $\Omega_{\rm mw}^{a}=2\pi\times 70~{}{\rm MHz}$ and $\Omega_{\textrm{mw}}^{b}=2\pi\times 200~{}{\rm MHz}$, the population of the target state can reach about $0.9916$ with a total evolution time $t_{\textrm{tot}}=0.3971~{}\mu{\rm s}$. Given that the independent evolution process of the core operation can reach perfection with $N=3$, the error caused by the preparation of the initial state and the return driving is approximately $0.0074$.
@@
Initial preparing and decoupling 0.0074
Spontaneous radiation 0.0006 0.0009 0.0010 0.0017
Atomic vibration (±$\pm 100~{}{\rm nm}$) 0.0009 0.0016 0.0020 0.0038
$P(N)_{O}$ 0.99 0.9821 0.9624 0.9165
As shown in Table. 3, we make an estimation on the population of complete states with $N=3\sim 6$ considering the experimental errors to make a comparison with the other protocol. In recent experiments, the fidelity of the two-qubit entanglement gate has been improved to $99.5\%$ [75, 76]. Based on this analysis, we can estimate the time and fidelity necessary to achieve complete graph states using CZ gates for comparison. For $N=3$, a minimum of three CZ gates are required, with an estimated fidelity of approximately $0.985$. For $N=6$, 15 CZ gates are needed, with an estimated fidelity of $0.928$, demonstrating performance comparable to our approach. However, for the gate-based approach, although a single gate operation time can be less than $300~{}{\rm ns}$, such CZ gate operations cannot be performed simultaneously. The evolution time will increase rapidly with increasing number of particles $N$. On the contrary, the evolution time required by our protocol is shorter, which can avoid some unknown errors. This reduction is crucial from the perspective of decoherence.
@@
In summary, we have introduced an efficient single-step technique for generating multiparticle complete graph states. Our approach relies solely on a constant nearest-neighbor interaction strength and a time-varying magnetic field which is optimized using the GRAPE algorithm. This methodology is applicable to various physical systems capable of implementing a spin-$1/2$ Heisenberg $XX$ chain, such as superconducting systems, trapped ion systems, and neutral atom systems. We have thoroughly investigated the protocol’s execution in the Rydberg atom array system, evaluating its resilience in the face of experimental errors. Our numerical findings underscore the robustness of our approach against challenges such as atomic position inaccuracies and pulse oscillations. Our one-step solution significantly reduces operating time while maintaining a relatively high estimated fidelity. We anticipate that this discovery will quickly contribute to the practical implementation of quantum computing and quantum error correction in neutral-atom systems.
Figure 8: Evolution governed by the Hamiltonian (20) with $N=3$. (a) The intensity of the constant amplitude pulse $B$ satisfies the relationship shown in Eq. (31). (b) Scanning results for the population of the target state with different parameters.
@@ acknowledgment
The authors thank Professor Re-Bing Wu of Tsinghua University for his guidance on the GRAPE optimal algorithm. The anonymous reviewers are also thanked for constructive comments that helped improve the quality of this article. This work is supported by the NFSC through Grant No. 12174048, the Fundamental Research Funds for the Central Universities (Grant No. 3122023QD26) and the Research Start-up Funds of the Shenyang Normal University (Grants No. BS202301, No. BS202303, No. BS202314).
@@ Appendix A A protocol for $N=3$ under constant amplitude magnetic field
$\psi_{1}(t)=\frac{e^{-\frac{3itB}{2}}}{2\sqrt{2}},$ (23)
$\psi_{2}(t)=\frac{e^{-i\frac{1}{2}t(4\sqrt{2}J+B)}[2+\sqrt{2}-(\sqrt{2}-2)e^{4i\sqrt{2}Jt}]}{8\sqrt{2}},$ (24)
$\psi_{3}(t)=\frac{e^{-i\frac{1}{2}t(4\sqrt{2}J+B)}[1+\sqrt{2}-(\sqrt{2}-1)e^{4i\sqrt{2}Jt}]}{4\sqrt{2}},$ (25)
$\psi_{4}(t)=\frac{1}{4}e^{\frac{itB}{2}}(\sqrt{2}\cos{\frac{4Jt}{\sqrt{2}}}-i\sin{\frac{4Jt}{\sqrt{2}}}),$ (26)
$\psi_{5}(t)=\frac{e^{-i\frac{1}{2}t(4\sqrt{2}J+B)}[2+\sqrt{2}-(\sqrt{2}-2)e^{4i\sqrt{2}Jt}]}{8\sqrt{2}},$ (27)
$\psi_{6}(t)=\frac{1}{4}e^{\frac{itB}{2}}(\sqrt{2}\cos{\frac{4Jt}{\sqrt{2}}}-2i\sin{\frac{4Jt}{\sqrt{2}}}),$ (28)
$\psi_{7}(t)=\frac{1}{4}e^{\frac{itB}{2}}(\sqrt{2}\cos{\frac{4Jt}{\sqrt{2}}}-i\sin{\frac{4Jt}{\sqrt{2}}}),$ (29)